OpenStudy (anonymous):

complete the identity (sin0+cos0)^2/1+ 2 sin0cos0

OpenStudy (anonymous):

(sinx +cosx)^2=sin^2 x+cos^2 x +2sinxcosx=1+2sinxcosx do you mean that?

OpenStudy (anonymous):

no the \[\theta\]

OpenStudy (anonymous):

so what is your question?

OpenStudy (anonymous):

all the zeros are \[\theta\]. and i need help with the equation

OpenStudy (anonymous):

he just changed the theta or θ into x =D

OpenStudy (anonymous):

yep..

OpenStudy (anonymous):

dang gurll thats a hard question theta kills me -__-